# and hierarchical hyperbolicity

### v2

@@ 4.2. Quotients by pseudo-Anosovs, low complexity
In this section we answer the question in the case where $U$ is a 4-punctured disk, whose mapping class group is the braid group on 4 strands, and therefore $\widehat{U}$ is a 5-punctured sphere $S_{5}$. As we will see, it is crucial for the proof that modding out powers of Dehn twists in the mapping class group of the five-punctured sphere yields a hyperbolic group. Later, we also discuss a different strategy to approach the general case.
@@
One can understand subgroups generated by powers of pseudo-Anosovs using the technology of rotating families; this is done in [DGO]. To prove the theorem, we will need an improvement of this technology from [dahmani:rotating]. We do not recall all the relevant definitions here, but we discuss the main points, giving the motivating example from [dahmani:rotating] along the way. A composite projection system is a set $\mathbb{Y}_{\ast}$ (for instance, the set of curves on a surface) together with additional data. One piece of data is, for each $X\in\mathbb{Y}_{\ast}$, a subset $Act(X)$ of $\mathbb{Y}_{\ast}$ (for instance, $Act(X)$ is the set of curves that intersect $X$). A composite rotating family on a composite projection system is given by groups $G_{X}$ of automorphisms of $\mathbb{Y}_{\ast}$ fixing $X$, for all $X\in\mathbb{Y}_{\ast}$, satisfying certain conditions (for instance, $G_{X}$ could be the subgroup generated by the Dehn twist around $X$).
We need the following technical proposition to make sure that the central extensions we will consider are still $\mathbb{Z}$-central extensions.
Proposition 4.4.
Let $G$ be a group acting by automorphisms on a composite projection system $\mathbb{Y}_{\ast}$, and let $Z$ be the kernel of the action. Suppose that we have subgroups $\{G_{X}\}_{X\in\mathbb{Y}_{\ast}}$ of $G$ such that the following hold.
  • $Z\cap G_{X}=\{1\}$.
  • $G_{gX}=gG_{X}g^{-1}$.
  • $G_{X}$ commutes with $G_{Y}$ if $X\notin Act(Y)$.
  • denoting $\hat{G}_{X}$ the image of $G_{X}$ in $G/Z$, we have that $\{\hat{G}_{X}\}$ forms a composite rotating family on $\mathbb{Y}_{\ast}$.
Then the quotient map $p\colon G\to G/Z$ restricts to an isomorphism
$N=\langle\hskip-1.99168pt\langle\{G_{X}\}\rangle\hskip-1.99168pt\rangle_{G}\to\langle\hskip-1.99168pt\langle\{\hat{G}_{X}\}\rangle\hskip-1.99168pt\rangle_{G/Z}=\hat{N}$.
Proof.
Clearly we have $p(N)=\hat{N}$, so we can restrict $p$ to $N$ to obtain a surjective homomorphism $\bar{p}\colon N\to\hat{N}$. In order to show that this is an isomorphism, we construct an inverse $\iota:\hat{N}\to N$, using the first presentation of $\hat{N}$ provided by [dahmani:rotating, Theorem 2.2], which applies by the fourth bullet. The $\hat{G}_{X}$ are generators for this presentation, and thus have natural maps to the $G_{X}$ by the first bullet. The relations for this presentation are conjugation relations, which are respected in view of the second bullet, and commutation relations, which are respected by the third bullet. Therefore, there is a well-defined homomorphism $N\to\hat{N}$, which when restricted to any $\hat{G}_{X}$ gives an inverse of $\bar{p}|_{G_{X}}$, and therefore $\iota$ is in fact the inverse of $\bar{p}$, as required. ∎
Proof of Theorem 4.3.
Throughout this this proof we always assume that $M$ is a multiple of some suitable $M_{0}\in\mathbb{N}_{>0}$.
First of all, we need $\mathcal{N}_{h}$ to intersect $\langle\tau_{\gamma}\rangle$ trivially, justifying that there is indeed a diagram as in the statement, with the vertical arrow on the left being an isomorphism. This follows from Proposition 4.4, with $G=B_{4}$, $Z=\langle\tau_{\gamma}\rangle$, and $G_{X}$ the conjugates of $\langle h^{M}\rangle$.
In order to show boundedness of the relevant central extension, we now argue that there is a further quotient central extension which is a central extension of a hyperbolic group. More precisely, we argue that there is a diagram
@@
where $\hat{H}$ is hyperbolic. Since the comparison map for hyperbolic groups is surjective [neumannreeves] (see also [mineyev]), by Lemma 2.12 we then get that the central extension from the statement is bounded, as required.
The required quotient $H$ can be taken to be the quotient of $G=B_{4}/\mathcal{N}_{h}$ by the normal subgroup $\mathcal{K}$ generated by suitable powers of all images of Dehn twists of $B_{4}$. We now explain why we can choose such a $\mathcal{K}$ that intersects trivially the image $Z$ of $\langle\tau_{\gamma}\rangle$ in $G$, and such that $G/\mathcal{K}Z$ is hyperbolic. This uses results from [short_HHG:I, short_HHG:II] about short HHG. We do not need the whole definition here; what is important is that these are HHG with a specified collection of cyclic subgroups, and if all of those are finite then the group is hyperbolic, see [short_HHG:I, Lemma 2.14]. All relevant short HHG below also satisfy the assumption called colourability from [DMS:stable, Definition 2.8], which from now on we omit. Modding out suitable powers of generators of any subcollection of the specified cyclic subgroups yields another short HHG by [short_HHG:II, Theorem 4.1], with associated collection of specified subgroups simply the images of the specified subgroups of the original short HHG.
First of all, $\hat{G}=\mathcal{MCG}(S_{5},p)/\mathcal{N}$ is a short HHG. This can be seen combining the following:
  • $\mathcal{MCG}(S_{5},p)$ is a short HHG, see [short_HHG:I, Subsection 2.3.1], with specified subgroups generated by Dehn twists.
  • We can add any cyclic subgroup generated by a pseudo-Anosov to the collection of specified subgroups of the short HHG structure of $\mathcal{MCG}(S_{5},p)$ by [short_HHG:I, Proposition 5.1].
  • The aforementioned [short_HHG:II, Theorem 4.1] allows us to conclude that modding out suitable powers of $h$ yields a short HHG (with the expected structure with specified subgroups coming from images of Dehn twists).
Now, for $\mathcal{K}$ as above we can consider its image $\hat{\mathcal{K}}$ in $\hat{G}$, and this is the kernel associated to a composite rotating family, see [short_HHG:II, Section 3.2]. This allows us to apply Proposition 4.4 and deduce that $\mathcal{K}$ intersects $Z$ trivially, as required. Moreover, $G/{\mathcal{K}}Z=\hat{G}/\hat{\mathcal{K}}$ is a short HHG by [short_HHG:II, Theorem 4.1], and all its specified subgroups are finite, which as mentioned above implies that the group is hyperbolic. This concludes the proof. ∎
@@
Under the assumption that “enough” hyperbolic groups are residually finite, [BHMS] constructs hyperbolic quotients of mapping class groups. One could in principle try to exploit these quotients, or variations, and a similar strategy as in the proof above to solve other cases of Question 4.2, or to solve it completely, but conditionally on the residual finiteness of the relevant hyperbolic groups. The main difficulty is ensuring that kernels to these hyperbolic quotients lift isomorphically to the corresponding central extension, as in Proposition 4.4.
4.3. Quotients by pseudo-Anosov, high complexity?
In the setting of Question 4.2, results from [DGO] allow us to write $\pi(\mathcal{N}_{h})$ as the directed union of groups $N_{i}$, each of which is a free product of finitely many conjugates of the cyclic group generated by the relevant power of $\hat{h}$. Each of the $N_{i}$ is convex-cocompact, and we expect that a variant of Theorem 2.19 allows one to extend any quasimorphism on $N_{i}$ to $\mathcal{MCG}(U)/\pi(\mathcal{N}_{h})$. If this extension could be realised in a way that the defect is uniformly controlled, then we would have a positive answer to the previous question, thanks to the following lemma:
Lemma 4.6.
Proof.
By using homogeneous representatives for one direction, and integer parts for the other, as we did before, we can reduce to the following statement. Let $\chi\colon N\to\mathbb{R}$ be a homogeneous quasimorphism. Suppose that $N$ can be expressed as a directed union of subgroup $\{N_{i}\}_{i\in I}$, and that for each $i\in I$ there is a homogeneous quasimorphism $X_{i}\colon G\to\mathbb{R}$ that extends $\chi|_{N_{i}}$, such that the defect of the $X_{i}$ is uniformly bounded. Then $\chi$ extends to a homogeneous quasimorphism $X\colon G\to\mathbb{R}$. Moreover, because $G$ is finitely generated, thus countable, we can assume that the directed set $I$ is just $\mathbb{N}$ with its well order.
$\alpha(g^{p})=\alpha\left(\prod\limits_{j=1}^{n}x_{j}^{p_{j}}\right)$
As pointed out in Remark 2.20, in our setting there is a single bounded cocycle that extends $\delta\chi_{i}|_{N_{i}}$. It could be possible to exploit this to find the extensions $X_{i}$ with uniformly bounded defect.

### v3

@@ 4.2. Quotients by pseudo-Anosovs
The first version of this paper answered Question 4.2 affirmatively in the case where $U$ is a 4-punctured disk, whose mapping class group is the braid group $B_{4}$ on four strands. However, since then, a complete answer has been obtained by Tao [tao2025extensiontheoremquasimorphisms, Corollary 1.7]. Since our proof was quite technical, instead of giving the details we shall highlight its overall strategy and some key ideas, since those are different from Tao’s and might apply in other settings.
@@
Proof idea.
Firstly, one proves that $\mathcal{N}_{h}$ intersects $\langle\tau_{\gamma}\rangle$ trivially, justifying that there is indeed a diagram as in the statement whose left vertical arrow is an isomorphism. This can be done using the machinery of rotating families, introduced in [DGO] to model the normal closure of a pseudo-Anosov (among many other examples).
Now, in order to show boundedness of the relevant central extension, it suffices to exhibit a further quotient central extension
@@
where $\hat{H}$ is hyperbolic. Indeed, the comparison map for hyperbolic groups is surjective [neumannreeves] (see also [mineyev]); hence, if the bottom extension exists, it is bounded, and in turn the central extension from the statement is also bounded by Lemma 2.12.
One can argue that $H$ can be taken to be the quotient of $G=B_{4}/\mathcal{N}_{h}$ by the normal subgroup $\mathcal{K}$ generated by suitable powers of all images of Dehn twists of $B_{4}$. In order to prove this, one needs techniques from [dahmani:rotating] to show that the intersection $\mathcal{K}\cap\mathbb{Z}$ is trivial, and then the machinery from [short_HHG:I, short_HHG:II], in particular [short_HHG:II, Theorem 4.1], to prove that $G/\mathbb{Z}\mathcal{K}$ is hyperbolic. Note that we can apply [short_HHG:I, short_HHG:II] because $\mathcal{MCG}(S_{5},p)$ is “sufficiently low-complexity” to be a so-called short HHG, as introduced in [short_HHG:I]. ∎
@@
Under the assumption that “enough” hyperbolic groups are residually finite, [BHMS] constructs hyperbolic quotients of mapping class groups. One could in principle try to exploit these quotients, or variations, and a similar strategy as in the proof above to solve some cases of Question 4.1, possibly conditionally on the residual finiteness of the relevant hyperbolic groups. The main difficulty is ensuring that kernels to these hyperbolic quotients, which morally arise as a union of kernels by rotating families, lift isomorphically to the corresponding central extension.